# Topological Peierls instabilities in more than one dimension

### v1

@@ Abstract
Introduction. The coupling between electron and lattice degrees of freedom is often key to understanding the emergence of electronic and structural order in complex materials. A textbook example is the Peierls transition, where a one-dimensional (1D) lattice becomes unstable toward dimerization, resulting in a reduced translational symmetry and an associated charge density wave [1]. This transition leads to the opening of a gap in the electron spectrum which lowers the system’s total energy, and stabilizes the distorted, insulating phase. Soon after periodic lattice distortions were observed in quasi-1D systems [2], charge density waves were discovered in effectively two-dimensional (2D) materials [3, 4]. Since then, understanding the ingredients introduced by the higher dimensionality has remained a topical question [5, 6, 7, 8, 9].
The fundamental connection between symmetries and topological phases in electronic systems [10, 11, 12, 13, 14] naturally raises the question of how electronic topology is altered when translational symmetry is reduced [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. This symmetry reduction leads to band folding into a smaller Brillouin zone, resulting in band crossings. While these crossings are generally subject to level repulsion and gap openings, symmetry constraints can enforce protected band crossings at specific wave vectors. Additionally, band-touching points may naturally arise in three-dimensional systems even in the absence of explicit symmetry protections [26, 27, 28]. Weyl and Dirac points serve as prominent examples of such crossings, where topological invariants protect the degeneracies and lead to robust surface states or exotic transport phenomena [29]. These crossings can prevent a full gap from developing at the Fermi level, leading to a topologically obstructed Peierls phase, in which levels are repelled from the Fermi energy, leaving behind a pseudogap in the spectrum.
As an example of this physics, we identify trigonal PtBi2, a noncentrosymmetric Weyl semimetal [30, 31, 32, 33, 34] that has recently garnered attention for its superconducting phase [35, 30, 31, 36, 37, 38]. Within a certain temperature range, this superconducting phase has been found to be exclusive to the topological Fermi arcs [31], underscoring the importance of a detailed understanding of the electronic structure in the normal phase.
In this Letter, we address the effects of translation symmetry reduction, focusing on systems with triangular lattice symmetry and connecting in a single framework the physics of honeycomb- and Kagome-based systems. We demonstrate different ways in which the electronic reconstruction associated with this symmetry reduction can stabilize a topological semimetallic phase. Finally, based on density-functional calculations of PtBi2, we show that this obstructed Peierls mechanism gives origin to the nontrivial topology of its normal phase.
@@
A 2D Peierls instability. Our starting point is a 2D triangular lattice. In Fig. 1, we present two examples of structural distortions that reduce its translational symmetry while preserving threefold rotational and reflection symmetries present in trigonal PtBi2. In these examples, sequences of short and long bonds can be identified, as in the 1D Peierls case, but the patterns can not be reduced to a mere stack of weakly coupled chains of dimerized bonds, underscoring their inherently 2D nature. In Fig. 1a), one out of three atoms is displaced out of the plane, resulting in a buckled, decorated honeycomb lattice. In Fig. 1b), one out of three atoms moves along a mirror-invariant line, leading to sequences of side-sharing small and large triangles along each of such lines.
@@
$F_{12}(\mathbf{k})$ $=e^{i\mathbf{k}\cdot(\mathbf{s_{1}}-\mathbf{s_{2}})}\Big{(}1+e^{-i\mathbf{k}\cdot\mathbf{a}}+e^{-i\mathbf{k}\cdot(\mathbf{a}+\mathbf{b})}\Big{)},$
@@
$D_{12}(\mathbf{k})=e^{i\mathbf{k}\cdot(\mathbf{s_{1}}-\mathbf{s_{2}})}\Big{(}1-e^{-i\mathbf{k}\cdot\mathbf{a}}-e^{-i\mathbf{k}\cdot(\mathbf{a}+\mathbf{b})}\Big{)}.$ (3)
@@
Figure 2: (a) Fragment of the distorted lattice [see Fig. 1b)] indicating the hopping amplitudes for short ($t+\delta$) and long ($t-\delta$) bonds. (b) Brillouin zones corresponding to the triangular lattice (dashed blue lines) and to the $\sqrt{3}\times\sqrt{3}$ supercell (black continous lines). The supercell Brillouin zone is rotated by $30^{\circ}$ and reduced in area by a factor of three relative to the original lattice. (c) Undistorted triangular lattice band structure folded into the supercell Brillouin zone. The reflection symmetry eigenvalues are indicated by $+$ and -- signs next to the bands. The linewidth indicates the unfolding weight which is $1$ for the lower band and 00 for the upper bands in this case. The Fermi level ($\varepsilon_{F}=0$) corresponds to an occupancy of a single electron per unit cell. (d,e) Band structure for the distorted case with (d) $\delta/t=-0.2$ and (e) $\delta/t=0.2$. (f) Density of states for $\delta/t=0.2$, 00, and $0.2$. Inset: Total electronic energy relative to the undistorted limit as a function of $\delta/t$.
@@
Thus, the sign of $\delta$ controls a phase transition between an insulator and a symmetry-enforced topological semimetal. Notably, the latter phase can be smoothly connected with Kagome-like systems, with a flat band and a Dirac cone at $K$ obtained in the present model in the $\delta\to t$ limit. Similar obstructions to a full gap opening can be obtained for the distortion in Fig. 1a). An important difference in this case is that the point symmetry of the atomic sites includes $C_{3}$. For this reason, the distorted electronic structure is smoothly connected to the honeycomb lattice case. Other possible distortions of the triangular lattice and the effects of longer range hoppings are analyzed in the SM [39].
@@
Figure 3: Band structure of a triangular AB stacked bilayer (a) without distortion and (b,c) with one of the layers dimerized: (b) $\delta/t=0.1$ and (c) $\delta/t=0.5$. Energies are measured with respect to the 1/3 filling chemical potential. As in Fig. 2, the linewidth indicates the unfolding weight and the reflection symmetry eigenvalues are indicated with $+$ and -- signs.
@@
[…]
@@
These nodes involve bands that at $\delta=0$ can be identified as folded and non-folded. They are not present for arbitrarily small distortions, rather, they emerge from the $\Gamma$-$M$ lines at a finite value of $\alpha\approx 0.39$ and persist for $\alpha$ up to $1$. They are monopoles of the Berry curvature which, without spin-orbit coupling, has as only sources orbital and site degrees for freedom.
The dimensionality, as well as the fact that the distortion not only reduces translational symmetry but also breaks inversion symmetry, are essential. Indeed, in a time-reversal symmetric three-dimensional system, codimension arguments indicate that this symmetry breaking can drive the system into a stable Weyl phase [27]. In PtBi2, while the Peierls distortion pushes spectral weight away from the Fermi energy, the created crossings introduce an obstruction–inextricably linked to the three-dimensional nature of the system– that prevents a full charge gap from opening. Importantly, as shown in Ref. [46], these crossings are robust to inclusion of the SOC, which essentially couples the nodes of each spin sector, leading to a splitting of the Weyl nodes in energy.
Conclusions. We have analyzed Peierls instabilities in systems with triangular lattice symmetry, where translation symmetry is reduced. As in the one-dimensional Peierls transition, the electronic reconstruction contributes to stabilizing the crystal distortion. However, the fate of the electronic structure is constrained by a topological obstruction that prevents the opening of a full gap. We have examined cases where this obstruction arises from a band crossing that carries a Berry phase of $\pi$, as well as cases involving a Chern number of 1. Our findings establish a smooth connection with the limiting cases of both Kagome and honeycomb lattices and suggest that materials based on distorted triangular lattices may host the electronic instability we describe. As an illustrative example, we have argued that the topology of trigonal PtBi2 can be effectively understood within this framework.
Acknowledgments. JF acknowledges useful discussions with Riccardo Vocaturo, Klaus Koepernik, Oleg Janson, Ion Cosma Fulga, and Jeroen van den Brink.

### v2

# Gapless topological Peierls-like instabilities in more than one dimension

@@ Abstract
Introduction. The coupling between electron and lattice degrees of freedom is often key to understanding the emergence of electronic and structural order in complex materials. A textbook example is the Peierls transition, where a one-dimensional (1D) lattice becomes unstable toward dimerization, resulting in a reduced translational symmetry [1]. This transition leads to the opening of a gap in the electron spectrum which lowers the system’s total energy, and stabilizes the distorted, insulating phase. Soon after periodic lattice distortions were observed in quasi-1D systems [2], they were also discovered in effectively two-dimensional (2D) materials [3, 4]. Since then, understanding the ingredients introduced by the higher dimensionality has remained a topical question [5, 6, 7, 8, 9].
The fundamental connection between symmetries and topological phases in electronic systems [10, 11, 12, 13, 14] naturally raises the question of how electronic topology is modified when translational symmetry is reduced. This issue has been considered in various settings, including how charge density waves affect gapless states on the surfaces of weak topological insulators [15], influence the bulk electronic structure of Weyl semimetals [16, 17, 18, 19, 20, 21, 22], and lead to the possible emergence of topological gaps [23, 24, 25].
The reduction of translational symmetry leads to band folding into a smaller Brillouin zone, resulting in band crossings that are generally susceptible to level repulsion and gap opening. While this can lead to a fully gapped insulating phase, the distorted system often remains metallic due to weak coupling between the lattice distortion and certain regions of the Fermi surface [26, 27, 28, 29, 30]. A natural question is whether the tendency toward gap opening can, in contrast, encounter obstructions rooted in topological properties rather than in weak coupling. Residual crystal symmetries can enforce protected degeneracies at specific wave vectors [31, 32, 33, 34, 35]. Moreover, band-touching points may arise naturally in three-dimensional systems even in the absence of explicit symmetry protection [36, 37, 38]. Weyl and Dirac points are prominent examples of such crossings, where topological invariants protect the degeneracies and give rise to robust surface states or exotic transport phenomena [39]. These features can obstruct the formation of a full gap at the Fermi level, resulting in a topologically obstructed Peierls phase, in which levels are repelled from the Fermi energy, leaving behind a pseudogap in the spectrum.
@@
In this Letter, we address the effects of translational symmetry reduction, focusing on systems with triangular lattice symmetry. We demonstrate distinct mechanisms by which the electronic reconstruction associated with this symmetry reduction can stabilize a topological semimetallic phase. Starting in two dimensions, we show that periodic lattice distortions that preserve point group symmetries can give rise to Dirac cones carrying a Berry phase, unifying within a single framework the physics of honeycomb- and Kagome-based systems. In three dimensions, we demonstrate that inversion-breaking periodic lattice distortions can stabilize a Weyl phase rather than an insulating one. As an example, we identify trigonal PtBi2, a noncentrosymmetric Weyl semimetal [40, 41, 42, 43, 44]. Based on density-functional theory calculations, we show that its normal phase reflects a Peierls-like mechanism obstructed by the emergence of Weyl nodes  [^1] We use the terminology of Peierls-like to emphasize that the main ingredient of the problems studied is the asymmetry in electron hopping amplitudes enabled by the breaking of translational symmetry, analogous to the 1D Peierls model..
2D Peierls instability. Our starting point is a 2D triangular lattice. In Fig. 1, we present two examples of structural distortions that reduce its translational symmetry while preserving threefold rotational and reflection symmetries present in trigonal PtBi2. In these examples, sequences of short and long bonds can be identified, as in the 1D Peierls case, but the patterns can not be reduced to a mere stack of weakly coupled chains of dimerized bonds, underscoring their inherently 2D nature. In Fig. 1a), one out of three atoms is displaced out of the plane, resulting in a buckled, decorated honeycomb lattice. In Fig. 1b), one out of three atoms moves along a mirror-invariant line, leading to sequences of side-sharing small and large triangles along each of such lines.
@@
$F_{12}(\mathbf{k})$ $=e^{-i\mathbf{k}\cdot(\mathbf{s_{1}}-\mathbf{s_{2}})}\Big(1+e^{-i\mathbf{k}\cdot\mathbf{a}}+e^{-i\mathbf{k}\cdot(\mathbf{a}+\mathbf{b})}\Big),$
@@
$D_{12}(\mathbf{k})=e^{-i\mathbf{k}\cdot(\mathbf{s_{1}}-\mathbf{s_{2}})}\Big(1-e^{-i\mathbf{k}\cdot\mathbf{a}}-e^{-i\mathbf{k}\cdot(\mathbf{a}+\mathbf{b})}\Big).$ (3)
@@
Figure 2: (a) Fragment of the distorted lattice [see Fig. 1b)] indicating the hopping amplitudes for short ($t+\delta$) and long ($t-\delta$) bonds. (b) Brillouin zones corresponding to the triangular lattice (dashed blue lines) and to the $\sqrt{3}\times\sqrt{3}$ supercell (black continous lines). The supercell Brillouin zone is rotated by $30^{\circ}$ and reduced in area by a factor of three relative to the original lattice. (c) Undistorted triangular lattice band structure folded into the supercell Brillouin zone. The reflection symmetry eigenvalues are indicated by ++ and - signs next to the bands. The linewidth indicates the unfolding weight which is 11 for the lower band and 0 for the upper bands in this case. The Fermi level ($\varepsilon_{F}=0$) corresponds to an occupancy of a single electron per unit cell. (d,e) Band structure for the distorted case with (d) $\delta/t=-0.2$ and (e) $\delta/t=0.2$. (f) Density of states for $\delta/t=0.2$, 0, and $0.2$. Inset: Total electronic energy relative to the undistorted limit as a function of $\delta/t$.
@@
Thus, the sign of $\delta$ controls a phase transition between an insulator and a symmetry-enforced topological semimetal. Notably, the latter phase can be smoothly connected with Kagome-like systems, with a flat band and a Dirac cone at $K$ obtained in the present model in the $\delta\to t$ limit. Similar obstructions to a full gap opening can be obtained for the distortion in Fig. 1a). An important difference in this case is that the point symmetry of the atomic sites includes $C_{3}$. For this reason, the distorted electronic structure is smoothly connected to the honeycomb lattice case. These findings place in perspective a common origin of Dirac cones in Kagome and honeycomb systems: a reduction of translational symmetry relative to a triangular lattice, while preserving a non-Abelian little group at the K points. In the SM, we present different distortions of a triangular lattice where a preserved point symmetry connecting sites of the enlarged unit cell stabilizes folding-induced band crossings [46].
@@
Figure 3: Band structure of a triangular AB stacked bilayer (a) without distortion and (b,c) with one of the layers dimerized: (b) $\delta/t=0.1$ and (c) $\delta/t=0.5$. Energies are measured with respect to the 1/3 filling chemical potential. As in Fig. 2, the linewidth indicates the unfolding weight and the reflection symmetry eigenvalues are indicated with ++ and - signs.
@@
We perform ab-initio calculations using the experimentally reported structure of noncentrosymmetric PtBi2 [51] ($S_{0}$) and of centrosymmetric Pt(Bi0.901Te0.099)2 [52] ($S_{1}$), in both cases representing the systems with the space group P31m [46]. This approach neglects effects of the Bi by Te substitution other than the change in the overall translation and inversion symmetries. We also consider intermediate structures obtained as $S_{\alpha}=(1-\alpha)S_{0}+\alpha S_{1},\,\,\,0\leq\alpha\leq 1$.
@@
Thus, as in the elementary models discussed earlier, the periodic lattice distortion in PtBi2 gives rise to two distinct electronic effects: a low-energy effect, consisting of a strong depletion of spectral weight near the Fermi energy, and a high-energy effect associated with an overall enhancement of hopping amplitudes. Both contribute to lowering the total electronic energy, which we find to be 50 meV lower in the distorted structure  [^2] For this comparison, we consider fully relaxed crystal structures..
@@
These nodes involve bands that at $\delta=0$ can be identified as folded and non-folded. They are not present for arbitrarily small distortions, rather, they emerge from the $\Gamma$-$M$ lines at a finite value of $\alpha\approx 0.39$ and persist for $\alpha$ up to $1$. They are monopoles of the Berry curvature which, without SOC, has as only sources orbital and site degrees for freedom [55, 56].
The case of PtBi2 shows how a Weyl phase can be stabilized by a periodic lattice distortion that breaks inversion symmetry, without relying on SOC. While conventional level repulsion is present between folded and non-folded bands, the emergence of stable crossing points at generic crystalline momenta becomes possible in three dimensions. This behaviour is rooted in a codimensionality argument: under broken inversion symmetry the crossing of two non-degenerate bands requires satisfying three constraints [37]. An external control parameter, in our case the amplitude of a periodic lattice distortion, enables tuning into such a gapless phase. Once the Weyl nodes are established, their stability is reflected in the fact that further changes to the control parameter shift their position in momentum space without opening a gap.
This stands in contrast to systems where charge density waves lead to insulating behavior [57, 58], or where metallicity is preserved due to weak coupling between low-energy states and the distortion potential [30]. Instead, PtBi2 falls into a strong-coupling regime, where significant spectral weight is removed from the Fermi level, yet a full gap does not open due to topological constraints. This strong level repulsion, together with charge conservation, helps explain why the low-energy electronic structure is relatively uncluttered compared to the centrosymmetric limit, with the Weyl nodes lying so close to the Fermi energy (Fig.4d). Importantly, these crossings remain robust upon inclusion of SOC[53], which acts to couple the nodes from each spin sector and split them in energy. Finally, since this mechanism does not rely on fine details such as lattice geometry or orbital character, it may be broadly relevant to other systems exhibiting inversion-breaking periodic lattice distortions in three dimensions.
Conclusions. We have analyzed Peierls-like instabilities in systems with triangular lattice symmetry, where translation symmetry is reduced. As in the one-dimensional Peierls transition, the electronic reconstruction contributes to stabilizing the crystal distortion. In higher dimensions, however, the fate of the electronic structure can be constrained by a topological obstruction that prevents the opening of a full gap. We have examined cases where this obstruction arises from a band crossing that carries a Berry phase of $\pi$, as well as a case involving a Chern number of 1. Our findings establish a smooth connection with the limiting cases of both Kagome and honeycomb lattices and suggest that materials based on distorted triangular lattices may host the electronic instability we describe. As an illustrative example, we have argued that the topology of trigonal PtBi2 can be effectively understood within this framework.
Acknowledgments. JF acknowledges useful discussions with Riccardo Vocaturo, Klaus Koepernik, Oleg Janson, Ion Cosma Fulga, and Jeroen van den Brink. Computational resources were provided by the HPC cluster of the Physics Department at Centro Atómico Bariloche (CNEA).